OpenStudy (anonymous):

What is 2688 in. reducded by 25%

OpenStudy (anonymous):

looks like 2016 in

OpenStudy (anonymous):

just take 3/4 of 2688

OpenStudy (unklerhaukus):

\[2688 in \times (1-25\%)\]\[=2688 in \times 75\%\]\[=2688 \times 0.75\]\[=2016\]

OpenStudy (anonymous):

take 2688*.25=672 then take 2688 and subtract the 672 for 2016